OpenStudy (anonymous):

what is the simplest form of 32^3/5?

ganeshie8 (ganeshie8):

write 32 as 2^5 then use the exponent rule \[\large (a^m)^n = a^{m\times n}\]

OpenStudy (anonymous):

what happens to the 3/5?

ganeshie8 (ganeshie8):

Lets see \[\huge \left(32\right)^{\frac{3}{5}} \] \[\huge \left(2^5\right)^{\frac{3}{5}} \] \[\huge 2^{5\times\frac{3}{5}} \] \[\huge 2^3\] \[\huge 8\]

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

yw!